A boat has 8 crew members,3 of whom can row only on one side and 2 only on the other side. Find the number of ways in which the crew can be arranged.

Answer by Edwin McCravy(20054)

Let the 8 crew members be

{L1,L2,L3,R1,R2,E1,E2,E3}

where the 3 L's can only row on the left side, the 2 R's
can only row on the right side, and the 3 E's can row on
either side.

Since all three L's must row on the left, we only need to choose
1 of the E's to row on the left.  

We can choose this E any of 3 ways.

The other 4 crew members will row on the right side. 

For each of those 3 ways to choose the fourth crew member for
the left side, there are 4! ways to arrange the 4 crew members on
the left side, and 4! ways to arrange the 4 crew members on the
right side.

Answer 3*4!*4! = 3*24*24 = 1728 ways.

Edwin